OpenStudy (anonymous):

please help I give medals What is the equation of the line that passes through (4, 3) and (-4, -1)?

OpenStudy (anonymous):

find the slope first by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (4,3) x1 = 4 and y1 = 3 (-4,-1) x2 = -4 and y2 = -1 now sub that into the slope formula and what do you get ?

OpenStudy (anonymous):

First find slope(m) : m=y2-y1/x2-x1 x1=4 x2=-4 y1=3 y2=-1 Means : m= -1-3/-4-4 Whats the slope(m) ?:) @Gronsniph

OpenStudy (anonymous):

ty bolth of you

OpenStudy (anonymous):

thats not the answer...that is just the slope

OpenStudy (anonymous):

do you still need help with the equation ?

OpenStudy (anonymous):

no I got it from here. you reminded me thanks

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

y = 2x + 7 right?

OpenStudy (anonymous):

No ! Your slope is 2 but i said you slope ! Again !